### v2

# TL;DR: Too Long, Do Re-weighting for Efficient LLM Reasoning Compression

Abstract
Large Language Models (LLMs) have recently achieved remarkable progress by leveraging Reinforcement Learning and extended Chain-of-Thought (CoT) techniques. However, the challenge of performing efficient language reasoning—especially during inference with extremely long outputs—has drawn increasing attention from the research community. In this work, we propose a dynamic ratio-based training pipeline that does not rely on sophisticated data annotations or interpolation between multiple models. We continuously balance the weights between the model’s System-1 and System-2 data to eliminate redundant reasoning processes while preserving the model’s reasoning capability. We validate our approach across models on DeepSeek-R1-Distill-7B and DeepSeek-R1-Distill-14B and on a diverse set of benchmarks with varying difficulty levels. Our method significantly reduces the number of output tokens by nearly 40% while maintaining the accuracy of the reasoning. Our code and data will be available soon.
Comparison of TLDR and baseline models in terms of average accuracy and token compression ratio. Higher values on both axes indicate better performance.
1 Introduction
“He that can have patience can have what he will.” — Benjamin Franklin
Recent efforts have developed reasoning-oriented Large Language Models (LLMs) capable of solving complex tasks. These models progressed from System 1 to System 2 paradigms (yu2024distilling21, ; li_system_2_reason, ). System 1 implementations, such as GPT-4o (GPT4o, ), LLaMA-3 (llama3, ), leverage rapid intuitive processing for immediate responses but struggle with complex reasoning tasks. In contrast, System 2 architectures such as DeepSeek-R1 (Deepseek-R1, ) are fine-tuned with extended thinking chains to promote deliberate analysis through iterative self-assessment, error mitigation, and verification. However, reasoning-oriented LLMs, which employ system-2 reasoning, tend to engage in excessive deliberation even for simple problems. This results in unnecessary exploration and planning, ultimately impairing their efficiency and practicality.
Direct approaches aim to addressing the issues of cognitive redundancy and excessive deliberation within reasoning LLMs. Training-free methods CoD ; yao2025activation ; TALE-EP include some that control the internal states of the model during reasoning through prompts or confidence-based techniques to compress the model. Alternatively, the pathway, exemplified by model merging, involves intervening in the parameters of the reasoning LLM to produce relatively concise solutions. Training-based methods primarily focus on sampling and synthesizing relatively concise reasoning paths on specified problem sets through various strategies TokenSkip ; TOPS ; CoT-Valve . These methods involve performing reinforcement learning SimPO ; thinkpruner ; O1-Pruner ; L1 or supervised fine-tuning (SFT) Overthink on reasoning LLMs, enabling the model to learn to generate more concise yet still correct reasoning paths.
Such methods typically require careful collection of problems and precise control of the data ratio for different lengths to achieve good results, leading to a complex process of parameter tuning and data construction. For example, TOPS TOPS requires pre-processing steps to manually label SFT data to construct length-sensitive models, while CoT-Valve CoT-Valve generates data by creating intermediate models through model interpolation for sampling. This construction process is often tedious TOPS , computationally expensive L1 , or difficult to control for quality CoT-Valve .
Demystifying Short/Long CoT Mixture in LLM Thinking Compression System-1 data (Short CoT on GSM8K-like easy problems) reduces reasoning redundancy on all problem levels. (Section 2) Short CoT reduces reasoning redundancy on simple questions and demonstrates generalization across varying levels of problem difficulty. System-2 data (Long CoT only on s1-like difficult problems) helps maintain performance. (Section 2) Incorporating a small proportion of Long CoT, particularly on challenging problems, can mitigate the accuracy degradation introduced by short CoT, while long CoT on simple questions doesn’t help much. Dynamic re-weighting of System-1/2 data builds effcient LLM Reasoning Compression. (Section 1) Driven by a simple intuition, we design a dynamic reweighting algorithm for system-1/2 data, achieving strong performance in LLM reasoning compression.
We investigate the effects of mixing short CoT and long CoT data on compressing reasoning. Our findings suggest that long CoT and short CoT induce divergent optimization directions in the model’s reasoning behavior. Increasing the proportion of short CoT encourages more concise reasoning patterns, but may lead to a decline in reasoning accuracy. In contrast, raising the proportion of long CoT helps preserve reasoning performance on complex tasks, though at the expense of reduced compression efficiency. This naturally raises the question: Can we identify an optimal Long-to-Short data mixture that strikes the best trade-off—maximizing reasoning efficiency while maintaining accuracy?
We base our approach on an intuitive motivation: when a model is thinking too long, it should re-weight more intuitive reasoning paths to simplify the thinking process. Conversely, when the thinking is too direct, it should incorporate more slow-thinking reasoning chains to encourage deeper contemplation. We propose a dynamic Thinking Length Data Re-Weighting method (TlDr), which dynamically balances the model’s complex reasoning using long CoT and efficient reasoning using short CoT data, enabling the model to eliminate redundant cognitive processes. First, we construct short CoT data for simple problems and long CoT data for complex problems. The model begins with an initial ratio and performs reasoning compression using mixed data. After completing a compression cycle, the model re-evaluates the expected benefits of System-1 CoT data and System-2 CoT data to achieve improved performance. Specifically, and in line with intuition, System-1 CoT data can enhance efficiency, so we use an efficiency metric to measure the expected benefit of System-1 data. System-2 CoT, on the other hand, improves reasoning accuracy, and we use an accuracy metric to measure the benefit of System-2 data in terms of reasoning capability.
Compared to various methods requiring fine-tuning data with different reasoning lengths, our approach enables dynamic ratio learning by utilizing the self-sampled long CoT data and the short CoT data constructed by the original instruct/base model. Through experiments on DeepSeek-Distill-7B/14B, our model achieves excellent compression results on the 7B/14B models, with only a slight decrease in reasoning capability.
2 Rethinking Short-Long CoT in Thinking Compression
We first constructed the short CoT data using simple problems and recorded how, as training steps increased, this subset contributed to token compression and accuracy retention across datasets of varying difficulty in math benchmarks.
Figure 1: Impact of Combining Short CoT and Long CoT in Fixed Ratios on Thinking Compression Performance and Token Cost. We assessed the variation decay rate in output token length and accuracy on datasets of various question difficulty, spanning from GSM8K to AIME. The Normalized Token/Acc metric detail please refer to Equ. 11 and Equ. 12.
We find that short CoT thinking data for simple problems (System-1 data) can help compress the token usage across questions of various difficulty levels. We leverage the short-cut solutions obtained from simple questions in GSM8K to fine-tune the model and then observe the token compression rates and accuracy drop rates across four datasets, ranging from simple to difficult: GSM8K, MATH500, AMC, and AIME. As shown in Figure 1, directly fine-tuning the long CoT model with short CoT data achieves good length compression for both simple and complex problems. We were pleasantly surprised to see that this form of length compression generalizes well across questions of all difficulty levels, and that it maintains strong performance on simple questions. However, this approach comes at a cost, as it leads to a significant decrease in reasoning ability on difficult problems. As this portion of the data is derived from intuitive CoT reasoning on simple problems, we denote it as System-1 data. It seems that directly using short CoT fine-tuning can only encourage the reasoning LLM to retain its System 1 reasoning abilities, while its ability for System 2 reasoning—slow and cautious thinking for complex problems—is largely lost.
We find that long CoT thinking data for difficult problems (System-2 data) can help maintain the model’s performance on challenging tasks, while simple question doesn’t help much. We sample with the s1 s1 like hard question prompt and then blend the System-2 data into the previous System-1 thinking dataset at a fixed short CoT vs. long CoT ratio: $0.8$:$0.2$. We then observe the token compression rates and accuracy drop rates across four datasets.
It is worth noting that, by contrast, when we mix more long CoT data from simpler questions, the model still experiences a significant drop in performance on difficult questions. Refer to the middle and bottom parts of Figure 1, where we mix the long CoT sampled from challenging problems with the short CoT from simple problems. As a baseline, we also mix long CoT and short CoT from simple problems. The long CoT from difficult problems achieves lower accuracy drop rates across different datasets while maintaining comparable token compression rates. We are unable to recover the original performance simply by using long CoT data from simple questions through data replay. Similar to the deliberate reasoning characteristic of the System-2 process on difficult problems, we refer to this part of the data as System-2 data.
A key question we directly address is whether a direct mixing ratio of the two types of data(System-1/2 data), can be employed for post-training the long CoT model, resulting in a solution that eliminates redundancy without compromising performance. Based on these observations, we propose a dynamic approach aimed at identifying the optimal Thinking Compression data.
Figure 2: Overview of TlDr: Starting with a System-2 model, we iteratively update it on both Short-CoT and Long-CoT samples. The ratios of both data sources are adjusted every several steps based on the current average model accuracy and token length from the validation set until convergence.
3 Thinking Length Dynamic Re-weighting
3.1 Short-Long CoT Reweighting with Relaxed Optimization
Algorithm 1 Long-to-Short (L2S) Dynamic Reweighting Pipeline
Domain data $\mathcal{D}_{\text{system-1}},\mathcal{D}_{\text{system-2}},\mathcal{D}_{\text{dev}}$; training steps $T$; batch size $b$; step size $\eta$; smoothing parameter $c\in[0,1]$ (e.g., $c=10^{-4}$ in our implementation)
Initialize proxy weights $\theta_{0}$
Initialize mixture weights $\alpha_{0}=(1/2,1/2)$
for $t=1$ to $T$ do
     Let $|x|$ denote token length of example $x$ (with $|x|\leq L$)
     Compute benefit of fine-tuning with System-1 data: $\lambda_{\text{system-1}}$ and System-2 data $\lambda_{\text{system-2}}$
     Update weights (entrywise exponential): $\alpha^{\prime}_{t}\leftarrow\alpha_{t-1}\cdot\exp(\eta\cdot\lambda_{t})$
     Renormalize and smooth: $\alpha_{t}\leftarrow(1-c)\frac{\alpha^{\prime}_{t}}{\sum_{i=1}^{k}\alpha^{\prime}_{t}[i]}+cu$
     Update proxy model weights $\theta_{t}$ using $L(\theta_{t-1},\alpha_{t})$ (e.g., via Adam, Adafactor)
end for
return $\frac{1}{T}\sum_{t=1}^{T}\alpha_{t}$
We formalize the thinking compression problem as an optimization task to determine the optimal ratio between System-1 and System-2 reasoning. We expect the model trained on mixed data to approach the superior performance of System-1 and System-2 in specific evaluation metrics. For model $M$ and input problem $x$, we define $T(y),C(y)$ as the token length and correctness of LLM output text $y$. We represent the System-1/2 ability bound as $\phi_{\text{sys-}i,\text{bound}}(x)$, in the following sections, we will abbreviate as $\phi_{\text{sys-}i,\text{bound}}(x)$
$\min_{\theta,\,\alpha\in(0,1)}\;L(\theta,\alpha)=\sum_{i=1}^{2}\alpha_{i}\cdot\delta_{i}$ (1)
$\delta_{i}=\phi_{\text{sys-}i,\text{bound}}(x)-\phi_{\text{sys-}i,\theta}(x)$ (2)
of which, $\phi_{sys-1,\theta}$ can be regarded as a metric for measuring the efficiency of the System-1 models. $\phi_{sys-2,\theta}$ can be regarded as an accuracy metric. In this way, the overall optimization objective is to minimize the gap between the model and the efficiency upper bound of System-1, as well as the reasoning capability upper bound of System-2, while simultaneously optimizing the model parameters to maximize both reasoning performance and efficiency.
$\phi_{sys-1,\text{bound}}$ $=-\mathbb{E}_{dev}[T(M_{s}(x))]$ (3)
$\phi_{sys-2,\text{bound}}$ $=\mathbb{E}_{dev}[C(M_{l}(x))]$ (4)
Setup for System-1/2 Mixed Data.
Since System-1 can provide fast and intuitive answers to simple problems, we use the short CoT model to modulate the data for the System-1 model. Since System-2 is designed to execute slow, logical reasoning for challenging problems, we employ the long CoT model to sample prompts from S1  s1 , retaining only the correct responses. Finally, we obtain $D_{system-1}=$<Simple Question, Short CoT> instruction pairs. For the harder problems within the System-1 domain, we used the long CoT model for sampling, resulting in a large amount of $D_{system-2}=$<Hard Question, Long CoT> instruction data.
3.2 Long-to-Short Data-Reweighting Tuning.
Step 1: Estimate the ideal upper bounds of efficiency and performance.
During training, we aim to continuously adjust the ratio of System-1 and System-2 data in the post-training phase, ensuring that the model retains the reasoning capabilities of the original long CoT model while achieving the efficiency of the short CoT model. Therefore, we set the accuracy upper bound, $\phi_{sys-2,\text{bound}}$, of the model obtained through mixed training to match the accuracy of the original long CoT model, while setting the token lower bound, $\phi_{sys-1,\text{bound}}$, of the mixed model to correspond to the data lower bound of the short CoT model we constructed.
$\phi_{\text{sys-2, bound}}$ $=\phi_{\text{sys-2},L}=\hat{\mathbb{E}}_{\text{dev}}[C^{L}(x)]=\frac{1}{K}\sum_{i=1}^{K}\mathbbm{1}[\text{Correct}(y_{i}^{L})]$ (5)
$\phi_{\text{sys-1, bound}}$ $=\phi_{\text{sys-1},short}=-\hat{\mathbb{E}}_{\text{dev}}[T^{S}(x)]=-\frac{1}{K}\sum_{i=1}^{K}\text{Token}(y_{i}^{S})$ (6)
Step 2: Thinking Compression Post-Train with dynamic System-1/2 reasoning weights
We dynamically evaluate the utility of System-1 and System-2 reasoning data during training, and, guided by the performance of a reference model, adjust the sampling ratio between the two data types in real time to optimize training effectiveness.
$\lambda_{\text{sys-1}}$ $=\max\left(\frac{\phi_{\text{sys-1, bound}}-\phi_{\text{sys-1},\theta_{\text{proxy}}}}{\phi_{\text{sys-1},\theta_{s}}-\phi_{\text{sys-1},\theta_{l}}},\ 0\right)$ (7)
$\lambda_{\text{sys-2}}$ $=\max\left(\frac{\phi_{\text{sys-2, bound}}-\phi_{\text{sys-2},\theta_{\text{proxy}}}}{\phi_{\text{sys-2},\theta_{l}}-\phi_{\text{sys-2},\theta_{s}}},\ 0\right)$ (8)
4 Experiments
Model Accuracy Generation Length A.C.R.
ASDiv GSM8K MATH AIME AMC Minerva Avg. ASDiv GSM8K MATH AIME AMC Minerva
7B Models
R1-Distill-Qwen 86.8 89.4 86.8 42.9 81.5 46.0 72.2 769 554 2861 6820 4510 3347
TALE-EP 80.4 89.1 84.3 40.0 80.0 42.3 69.3 509 450 1994 6520 3892 2242 22.3%
ConciseCoT 86.0 89.5 86.2 41.7 79.6 46.0 71.5 532 457 2330 6587 4245 3347 12.7%
Avg. Merging 92.8 70.1 58.6 0.05 39.6 29.8 48.4 622 8552 8540 8501 8542 8544 3.2%
Task-Arithmetic-Merging 83.3 84.6 74.6 20.0 63.5 39.6 61.0 321 383 907 2500 1311 794 61.3%
Ties-Merging 74.4 69.7 59.8 13.6 42.5 23.2 47.2 1114 2475 4086 6767 5195 4306 0.1%
Ties-Dare-Merging 75.9 72.3 65.4 14.6 45.6 24.3 49.6 1036 2073 2934 5483 3698 2938 8.3%
Overthink 86.6 89.6 87.2 38.7 79.6 45.2 71.1 773 555 2898 6766 4558 3407 0.1%
ThinkPrune 90.6 92.1 91.0 43.3 86.2 45.6 74.8 653 587 2379 6207 3739 2762 12.6%
[…]
$\Delta$ +6.2 -1.7 +0.6 -1.7 +1.6 -1.5 +0.7 -622 -301 -1305 -452 -1124 -1896
14B Models
R1-Distill-Qwen 80.5 92.5 86.4 43.4 79.6 48.2 71.7 476 679 2951 6701 4584 3270
TALE-EP 77.5 92.4 85.4 49.2 80.3 50.0 72.5 369 555 2248 6551 4179 2731 15.4%
ConciseCoT 74.0 92.4 85.6 51.6 82.3 47.1 72.2 369 555 2066 6267 3878 2605 18.8%
Avg. Merging 94.8 90.3 73.0 10.8 55.0 44.1 61.3 167 366 5158 6364 5668 1084 30.5%
Task-Arithmetic-Merging 86.5 86.5 74.2 13.3 55.3 36.0 58.6 238 368 870 2813 1411 1050 60.2%
Ties-Merging 79.6 91.3 82.6 25.4 72.5 37.1 64.8 242 542 1919 5913 3158 1850 31.8%
Ties-Dare-Merging 80.7 91.8 84.8 25.4 75.3 34.9 65.4 274 467 1870 5747 3182 1877 33.0%
Overthink 79.3 92.3 88.0 45.8 82.8 45.6 72.3 451 679 2893 6700 4464 3715 1.6%
ThinkPrune 80.6 93.7 89.0 50.8 88.7 50.7 75.6 379 563 2177 5778 3327 2234 22.8%
[…]
$\Delta$ +8.0 -1.6 +0.2 -0.1 +4.2 +0.5 +2.1 -318 -439 -859 -298 -745 -1093
Table 1: Performance comparison of TlDr with baselines. The accuracy is measured by sampling multiple responses from the LLMs and taking the average to reduce variance. * denotes the CoT-Valve CoT-Valve result that we reproduced using the official dataset. $\Delta$ refers to TlDr in comparison with Original. Math’ and Minerva’ refer to MATH500 and MinervaMath datasets, respectively. A.C.R. means the token compression ratio computed by Eq. 9. In the table: yellow represents prompt-based methods; green highlights Merging-based methods; red indicates SFT-based and RL-based methods.
4.1 Experimental Setup
Model Accuracy Generation Length
GSM8K MATH AIME AMC Minerva Avg. GSM8K MATH AIME AMC Minerva Avg.
7B Models
Original Model 89.4 86.8 42.9 81.5 46.0 69.3 554 2861 6820 4510 3347 3618
[…]
-Static-Mixture 87.1 84.8 39.7 73.1 35.5 64.0 236 1221 5322 2560 1544 2177
-TlDr 87.7 87.4 41.2 83.1 41.0 68.1 253 1556 6368 3386 1434 2599
Table 2: Performance comparison of TlDr with static baselines. The accuracy is measured by sampling multiple responses from the LLMs and taking the average to reduce variance. & denotes the CoT-Valve CoT-Valve result that we reproduced using the official dataset. Math and Minerva mean MATH500 and MinervaMath datasets.
Model Accuracy Generation Length
GSM8K MATH AIME AMC Avg. GSM8K MATH AIME AMC Avg.
DeepSeek-R1-Distill-Qwen-7B
Original Model 89.4 86.8 42.9 81.5 75.2 554 2861 6820 4510 3686
[…]
-L1-same 86.4 88.6 42.2 84.6 75.4 301 2301 5875 3784 3056
-L1-lower 86.4 87.6 45.1 84.6 75.9 312 1831 5675 3807 2906
-L1-higher 86.1 88.4 45.5 83.3 75.8 292 2589 6007 3746 3158
Table 3: Performance comparison of TlDr with budget-aware baseline, L1 (L1, ). The accuracy is measured by sampling multiple responses from the LLMs and to reduce variance. The terms same, lower, and higher refer to setting the budget to match our results, 20% lower, and 20% higher, respectively. MATH means MATH500 dataset.
Model Accuracy Generation Length
GSM8K MATH AIME AMC Avg. GSM8K MATH AIME AMC Avg.
DeepSeek-R1-Distill-Qwen-7B System-1 Short CoT Ablation
Original Model 89.4 86.8 42.9 81.5 60.1 554 2861 6820 4510 2949
-TlDr-Easy 87.7 87.4 41.2 83.1 59.9 253 1556 6368 3386 2313
-TlDr-Medium 88.2 86.2 41.5 31.3 61.8 318 2083 6604 3945 3238
-TlDr-Hard 83.6 80.2 30.0 65.3 64.8 495 2970 6874 4947 3822
DeepSeek-R1-Distill-Qwen-7B System-2 Long CoT Ablation
Original Model 89.4 86.8 42.9 81.5 60.1 554 2861 6820 4510 2949
-TlDr-Easy 83.9 86.8 42.5 83.4 74.2 446 2639 6580 4047 3428
-TlDr-Medium 91.6 87.6 40.4 81.5 75.3 542 2761 6553 4116 2950
-TlDr-Hard 87.7 87.4 41.2 83.1 74.8 253 1556 6368 3386 2828
Table 4: An ablation study on the difficulty levels of ShortCoT and LongCoT was conducted during the construction of the ShortCoT and LongCoT dataset. The accuracy is measured by sampling multiple responses from the LLMs and taking the average to reduce variance. & denotes the CoT-Valve CoT-Valve result that we reproduced using the officially dataset. MATH means MATH500 dataset.
Figure 3: Comparison of accuracy and generation length between Vanilla CoT and our TlDr method on four benchmark datasets (GSM8K, MATH500, AIME, AMC) using DeepSeek-R1-Distill-Qwen models. TlDr consistently reduces generation length while maintaining or improving accuracy across both 7B and 14B model scales.
Figure 4: Frequency comparison of different keywords. The figure illustrates the distribution of exploratory, checking, and reflective keywords across datasets. Exploratory Keywords: wait, Reflective Word: but, Checking Words: make sure/confirm/verify/check, TlDr significantly reduces the presence of such words, reflecting its ability to produce streamlined and efficient reasoning steps.
Datasets and Metrics.
Following prior efforts, we evaluate TlDr on several widely-used benchmarks that span a broad range of difficulty levels, including ASDiv miao2021diverse , GSM8K (cobbe2021training, ), MATH-500 (hendrycks2021measuring, ), and AIME2024 (AIME2024, ), AMC (AMC2023, ) in Table 1. To ensure the stability of the evaluation, we performed multiple samplings for each dataset and took the average accuracy. For GSM8K, MATH500, and GPQA, we sampled each question 4 times and took the average accuracy of the 4 solutions. For AIME24 and AMC23, we sampled each solution 8 times and took the average accuracy of the 8 solutions. The token count was calculated using the corresponding tokenizer of the language model, and our evaluation environment was modified using the Skythought library.
Baselines.
We compared three types of baselines:
Prompt-based methods. We compared our approach with the well-known prompt-based baselines in the community, including TALE-EP TALE-EP , which requires the prompt to be as simple as possible, and ConciseCoT ConciseCoT , which demands the use of the most concise CoT steps during step-by-step reasoning.
Model-Merging-based Methods. Model Merging leverages the rich knowledge from short CoT Instruct and the long CoT model for model fusion, aiming to achieve the shortest yet most effective reasoning process. We compared this approach with the Avg. Merging method used in Kimi-1.5 team2025kimi ; Merging-Long2Short and some advanced merging method, like Task-Arithmetic-Merging, Ties-Merging, Ties-Dare-Merging, discussed in Merging-Long2Short .
Reward-based Methods. ThinkPruner thinkerprune uses progressive compression of RL training length to improve the effectiveness of context utilization during exploration. $SimPO_{shortest}$ was introduced in Overthink kumar2025overthink to adjust the effectiveness of the RL algorithm by length-guided RL training.
We find that our approach can maintain accuracy across datasets of varying difficulty while achieving satisfactory compression rates. In contrast, merging-based methods still suffer from significant performance drops on challenging problems. Compared to CoT-Valve and ThinkPrune, our method attains excellent compression rates, particularly on the ASDiv and GSM8K datasets, where over-exploration tends to occur. CoT-Valve, as an SFT-based approach, requires carefully designed model blending and the construction of length-diverse datasets for dynamic learning. In contrast, our method only requires straightforward data sampling and adaptive mixing ratios, achieving adaptive reasoning in a much simpler way.
4.2 Comparison of Dynamic Compression vs. Static Compression Data.
We compared our dynamic algorithm with several simple static data mixing methods, including directly mixing data according to a simple ratio and constructing length-uniform data using parameter interpolation. MixChain-Z-GSM8K is a Long2Short dataset proposed by CoT-Valve in Table 2.
4.3 Ablation of Different System-1/2 Source
We also discovered that incorporating higher-difficulty CoT data into a short-long mixed dataset could effectively eliminate redundancies in CoT for their compressed version. However, direct mixing could lead to performance degradation. After introducing a dynamic ratio method, we found that flexibly adjusting the ratio could effectively maintain performance in Table 4. We categorized the sources of questions in the thinking compression data into three difficulty levels: easy, medium, and hard. easy questions are from GSM8K, medium questions are from the training set of MATH500, and hard questions are from the s1 prompt questions.
Short CoT Compression Generalization Analysis of Easy-to-Hard. We tested the construction of System-1 data, examining the composition of data from different thinking compression sources. Our experiments found that constructing data based on low-difficulty problems could significantly reduce the token count of high-difficulty problems while maintaining performance. We found that using lower-difficulty problems to construct thinking compression data for redundancy removal can further generalize to higher-difficulty problems.
Long CoT Performance Generalization Analysis of Hard-to-Easy We also conducted an analysis of the following aspects: during the sampling of long CoT, we utilized data from three distinct sources—easy, medium, hard prompt. Our findings reveal that only by constructing long CoT using hard problems and dynamically adjusting their proportions during training can we recover the original performance associated with long CoT. This strategy effectively mitigates the risk of forgetting in reasoning capabilities during continual learning.
4.4 Comparison with Token Budgeted-Aware Model
We compared our redundancy reduction method with both quota-controlled models and reasoning models under the same token budget, in order to evaluate the effectiveness of our approach relative to explicit quota-based control. The results show that our method achieves higher reasoning accuracy than both the L1 L1 baselines under the same token quota. Furthermore, our approach demonstrates more efficient utilization of context length and does not require explicitly specifying a reasoning quota, offering a more flexible and adaptive inference mechanism. TlDr demonstrates stronger compression efficiency on simple problems.
4.5 Analysis of Thinking Patterns: Reflections & Solutions
We compared our method with other thinking compression methods in terms of their impact on changes in cognitive patterns  RedStar of the solution in Figure 3. We performed fine-grained statistical analysis on the results across different datasets and various levels of MATH500. Our analysis demonstrates that our approach effectively compresses the internal redundancy and reflects the properties of the solution patterns. TlDr effectively reduces the reliance on such macro reasoning patterns in benchmarks like GSM8K and MATH500, thereby avoiding excessive allocation of computational budget. Notably, for more challenging problems, the model still retains a significant degree of macro reasoning behavior to preserve its System-2 reasoning capabilities.
5 Related Work
Efficient System-2 Reasoning.
Despite the strong generalization and reasoning abilities demonstrated by the system 2 reasoning paradigm, the auto-regressive nature of LLMs imposes a significant reasoning burden, limiting their application in agent-based or edge scenarios OS-ATLAS . To address this, various approaches have emerged to improve the reasoning efficiency. These methods can be broadly categorized into two types. One category focuses on building adaptive reasoning-budget. Within this, some training-free methods like CoD CoD and TALE-EP TALE-EP impose budget constraints to control overall reasoning cost. Budget-sensitive models such as L1 L1 , TOPS TOPS , o1 Pruner o1_pruner , and K1.5 team2025kimi add length penalties during the post-training. Some work CoT-Valve ; jiang2025drp ; wang2024explore ; chain-of-reasoning synthesizes diverse-length CoT data, while TOPS TOPS samples budget-sensitive versions using a data model, and C3oT C3oT compresses original LLM output to train jointly on shorter CoTs. Other approaches involve external models or switching mechanisms Reasoning-in-Flux to allocate budgets. Routellm RouteLLM uses multiple routers to find the most suitable reasoning model, while Self-REF Self-REF employs confidence scores to route based on reasoning difficulty. Another category concentrates on building efficient representations. TokenSkip TokenSkip selects data based on token importance for compressed reasoning and more concise thought chains. COCONUT COCONUT explores more efficient reasoning in the latent space. ICoT-KD CCoT and CCoT CCoT attempt to build more efficient reasoning strategies in the hidden space, while Token Assorted combines hidden space and text-based reasoning to balance interpretability and efficiency. Heima Heima extends hidden-space reasoning to multi-modal models. Unlike prior efforts, TlDr compresses the length of the reasoning chain without introducing complex data construction and without compromising the original reasoning LLM’s inferential representations.
Data Balancing in Pre/Post Training.
The quality and proportion of data are critical during both the pre-training and post-training phases. In the pre-training stage, data quality and proportion are primarily managed through filtering and reweighting. Pre-training data filtering, extensively studied to boost model performance and training efficiency liu2024datasets ; albalak2024survey , typically involves steps like language filtering bigscience ; chowdhery2022palm , quality filtering raffel2023exploring ; rae2022scaling , content filtering xu2021detoxifying ; longpre2023pretrainers , and deduplication hernandez2022scaling ; lee2022deduplicating . While these methods significantly enhance corpus quality, their static nature can hinder dynamic adjustments during training, potentially discarding valuable data muennighoff2023scaling and introducing biases gururangan-etal-2022-whose ; longpre2023pretrainers ; dodge2021documenting . Similarly, in the post-training stage, an appropriate proportion of data with varying characteristics is crucial for optimizing final performance bpo ; zhang_balancing . For example, DeepMath-103K generates a large volume of data with evenly distributed difficulty for training DeepMath-103K , SRPO designs a dynamic sampling approach to filter out samples that are consistently answered correctly, thereby improving inference efficiency SPRO . EffiCode (huang2024effi, ) develop a self-optimization process based on overhead profiling and iterative refinement, resulting in a high-quality dataset for fine-tuning LLMs to produce more efficient and accurate code. To the best of our knowledge, we are the first to introduce a re-weighting mechanism into thinking compression. By employing simple strategies to construct short and long CoT, we enable the model to dynamically compress its reasoning process.
6 Conclusion
This paper introduces TlDr, an innovative method designed to compress the reasoning processes of LLMs without sacrificing accuracy. By dynamically re-weighting the influence of system 1 (concise reasoning) and system 2 (detailed reasoning) data during the training process, TlDr allows LLMs to eliminate unnecessary steps for simpler problems while still engaging in deep contemplation for complex tasks. TlDr avoids the laborious data collection and hyperparameter tuning typically required by other compression methods, offering a more practical solution for developing LLMs that are both efficient and accurate.
Appendix A Appendix
A.1 Training Details
Due to the need to evaluate accuracy and token count on a validation set every n steps, our validation set consists of 512 questions sampled from past questions in AIME-1983 to AIME-2023. The original ratio for shortcot and longcot is set to 0.5:0.5, with an evaluation interval of every 32 steps. The model is allowed to train for a total of 2000 steps, and the learning rate is set as a constant at 1e-5. For the 7B and 14B models, we conducted training on two 8-GPU (80GB) machines, with one 8-GPU machine performing vllm inference and the other performing training. Every n steps, parameter synchronization is executed using vllm’s parameter sync function. For the 32B model, we performed SFT on five 8-GPU nodes and deployed the inference service on one node.
We ultimately select the checkpoint with the shortest token length among those whose accuracy on the validation set is no less than 30% of that achieved by the original long CoT.
A.2 Metrics
A.2.1 Compression Rate
We provide more details on the compression rate in the main table, where the compression rate is defined as:
[…]
A.2.2 Normalized Metric
We report two normalized metrics to facilitate fair comparisons: Normalized Accuracy and Normalized Token Length. They are defined as follows:
Normalized $\text{Normalized Accuracy}=\frac{\#\text{Acc}_{\text{current}}}{\#\text{Acc}_{\text{original}}}$ (11)
Normalized $\text{Normalized Token}=\frac{\#\text{Token}_{\text{current}}}{\#\text{Token}_{\text{original}}}$ (12)
A.3 Data Construction Detail
For long CoT, we use the prompt from dataset s1.1  [37]. Each sample is generated 8 times using the original model. For short CoT, to avoid inconsistencies in the system prompt format, we adopt the short CoT construction method from AdaR1. We annotate 10 randomly selected questions from GSM8K using the instruct model, then fine-tune the long CoT model to overfit on them. For the GSM8K training set, we sample and retain only the examples with correct answers.
A.4 Evaluation Detail
We use the DeepSeek-R1-Distill model and apply a temperature setting of 0.7, which is the primary recommendation in QwQ-Preview, for evaluating all models. All datasets are restricted to an 8K context window for output generation. Meanwhile, considering the relatively small sizes of the AMC and AIME datasets, we sample 8 responses per question and compute the average.
A.5 Evaluation Framework
We use skythought-eval [^1] https://github.com/NovaSky-AI/SkyThought as the framework, which supports accelerating long CoT reasoning evaluation with vLLM. The version of vLLM we use is 0.6.3.
A.6 Evaluation Dataset Detail
We provide an overview of all datasets used in the following sections.
  • ASDiv: A diverse simple English math word problem corpus for evaluating the capability of various MWP solvers. It contains 2,305 MWPs that cover more text patterns and most problem types taught in elementary school.
  • GSM8K: A high-quality benchmark comprising 8,500 human-written grade school math word problems that require multi-step reasoning and basic arithmetic, each labeled with a natural language solution and verified answer. The 1,319-question test set emphasizes sequential reasoning and is primarily solvable by upper-grade elementary school students.
  • MATH500: a challenging benchmark of 500 high school competition-level problems spanning seven subjects, including Algebra, Geometry, Number Theory, and Precalculus. Each problem is presented in natural language with LaTeX-formatted notation, offering a strong measure of mathematical reasoning and generalization across diverse topics.
  • AIME2024: a dataset containing 30 problems from the 2024 American Invitational Mathematics Examination (AIME), a prestigious high school mathematics competition for top-performing students. Each problem is designed to require deep mathematical insight, multi-step reasoning, and precise problem-solving skills.
  • AMC: The AMC dataset consists of all 83 problems from AMC12 2022 and AMC12 2023, extracted from the AoPS wiki page. We used a subset of this data containing 40 problems.
  • MinervaMath: MinervaMath is a high-difficulty math problem dataset containing 272 challenging problems.
A.7 Reproduce Details
ConciseCoT & TALE-EP For the prompt-based baseline, we list the prompts used in Prompt 5.
OverThink For the MATH12K dataset, we sample each problem 8 times. The shortest correct sample is selected as the chosen sample, and the longest sample is selected as the rejected sample. The model is trained for 1 epoch.
ThinkPruner In our reproduction, we use the competition-level training data provided in the original paper and train the model for 10 epochs with a learning rate of 1e-6. The maximum response length is set to 4096 tokens. We follow their early stopping strategy to select the optimal checkpoint for evaluation.
CoT-Valve Since CoT-Valve does not report performance on all datasets, we reproduced the results using the public datasets released by CoT-Valve. We followed the training settings officially reported in the paper, using LoRA=2 to fine-tune all models. The dataset version used is Mix-Chain-Z-GSM8K. All models were fine-tuned for 5 epochs on 8 GPUs with 80GB of memory each.
Evaluation Prompt on Dataset === EVALUATION PROMPT FOR GSM8K === <|begin_of_sentence|>Please reason step by step, and put your final answer within ’boxed’. <|User|>query<|Assistant|>Given the following problem, reason and give a final answer to the problem. Problem: {question} Your response should end with T̈he final answer is [answer]ẅhere [answer] is the response to the problem. <think> === EVALUATION PROMPT FOR MATH500 === <|begin_of_sentence|>Please reason step by step, and put your final answer within ’ boxed’. <|User|>{query}<|Assistant|>Return your final response within ’boxed’. {problem}. <think> === EVALUATION PROMPT FOR AIME24 === <|begin_of_sentence|>Please reason step by step, and put your final answer within ’ boxed’. <|User|>query<|Assistant|>Return your final response within ’ boxed’. {problem}. <think> === EVALUATION PROMPT FOR AMC === <|begin_of_sentence|>Please reason step by step, and put your final answer within ’boxed’. <|User|>query<|Assistant|>Return your final response within ’boxed’. {problem}. <think> === EVALUATION PROMPT FOR MINERVAMATH === <|begin_of_sentence|>Please reason step by step, and put your final answer within ’boxed’. <|User|>query<|Assistant|>Return your final response within ’boxed’. problem. <think>
Figure 5: Evaluation Prompt for GSM8K, MATH500, AIME24, AMC, MinervaMath
L1 In L1 reproduction on the 7B System-2 model, we utilize the L1-Exact reward function and limit the token length to between 100 and 4,096 tokens, while setting the token difference penalization parameter $\alpha$ to 0.0003, as described in the paper. We follow their original prompt by appending "Think for $n_{token}$ tokens" to the end of the question. In inference, the token budget is set to the same number as the average tokens from our method across the evaluated benchmarks.
Appendix B Case Study
To better understand the behavioral differences between baseline and TLDR strategies, we conduct a qualitative analysis using the DeepSeek-R1-Distill-Qwen-7B model. Case studies are drawn from three representative math datasets: GSM8K, AIME, and MATH500. As shown in Figures 68, the baseline model tends to generate verbose reasoning paths with redundant or speculative content. In contrast, TLDR produces significantly more concise outputs while maintaining correctness and logical structure. These examples demonstrate TLDR’s ability to suppress unnecessary reasoning tokens—such as exploratory or reflective phrases—leading to more efficient and focused reasoning processes.
Figure 6: Comparison of Reasoning process on GSM8K: Baseline vs. TLDR.
Figure 7: Comparison of Reasoning process on AIME: Baseline vs. TLDR.
Figure 8: Comparison of Reasoning process on MATH500: Baseline vs. TLDR.